OpenStudy (anonymous):

15(2 - x) = 13(3 - x)

OpenStudy (lilg132):

30 - 15x = 39 - 13x 13x - 15x = 39 - 30 -2x = 9 x = -4.5

OpenStudy (anonymous):

and in fraction form?

OpenStudy (lilg132):

-9/2

OpenStudy (anonymous):

\[\frac{-9}{2}\]

OpenStudy (anonymous):

ughhh why cant i get this.... Thanks i'll try and get the next one by myself ;)

OpenStudy (anonymous):

\[-\frac{9}{2} \]